# Complex structures of the Gibbons-Hawking ansatz with infinite topological type

### v1

@@ Abstract
In this paper, we study the complex structures of complete hyperkähler four-manifolds of infinite topological type arising from the Gibbons-Hawking ansatz. Under a natural genericity assumption on the monopole configuration, we show that the resulting manifold is biholomorphic to a hypersurface in $\mathbb{C}^{3}$ defined by an explicit entire function, and we prove that this description persists for almost all complex structures in the associated $S^{2}$-family of compatible complex structures. When the genericity condition is relaxed, we further demonstrate that the manifold is biholomorphic to the minimal resolution of a singular surface. Thus, we partially extend LeBrun’s celebrated work [18] to the context of countably many monopoles.
@@ 1 Introduction
In their pioneering work, Gibbons and Hawking [9], [10] introduced a construction of four-dimensional hyperkähler metrics with a tri-Hamiltonian $S^{1}$-action, now known as the Gibbons-Hawking ansatz. Hitchin [14], through the twistor approach and the study of monopole moduli spaces, developed a coherent picture of the hyperkähler geometry behind these constructions. Kronheimer [17] described asymptotically locally Euclidean spaces arising as hyperkähler quotients. Bielawski [3] analyzed complete hyperkähler manifolds with local tri-Hamiltonian actions, placing a wide range of examples into a unified framework. In these works, the relevant configurations of monopole points are finite, and the resulting manifolds are of finite topological type. On the other hand, Anderson, Kronheimer and LeBrun [1] used Gibbons-Hawking metrics to construct complete Ricci-flat hyperkähler manifolds of infinite topological type. Goto [11] later gave a different viewpoint in terms of hyperkähler quotients, closely related to Kronheimer’s construction [17] of hyperkähler manifolds of type $A_{k}$. Subsequent work by Hattori [12] studied the volume growth and asymptotic geometry of these type $A_{\infty}$ manifolds. Closely related constructions appear in the theory of hypertoric manifolds, where Dancer and Swann analyzed some examples in [5, 6].
In this manuscript, inspired by the works of [1, 16, 18, 4], we investigate the complex structures of complete hyperkähler manifolds $M$ of infinite topological type obtained from the Gibbons-Hawking ansatz. Our first result (Theorem 1.1) gives an existence criterion for the Gibbons-Hawking potentials. The corresponding existence statement was already proved in work of Kalafat and Sawon [16] by an argument based on the maximum principle. We recover the criterion by a different method, using potential theory and the language of Riesz measures. Based on the method in LeBrun [18], we show that there exists a complex structure in the hyperkähler family on $M$ under which $M$ is biholomorphic to a hypersurface in $\mathbb{C}^{3}$ under a generic assumption (Theorem 1.3). LeBrun’s work treats the case where the circle action has finite fixed points and our description extends his picture to the case of a countable configuration of monopoles. After relaxing the genericity assumption, we further prove that the manifold can be realized as the minimal resolution of a singular hypersurface of a similar form (Theorem 1.4). In this way we relate the complex structures arising from the Gibbons–Hawking ansatz with infinite topological type directly to the geometry of the monopole set.
@@ 1.1 Background of Gibbons-Hawking ansatz
Let $U\subset\mathbb{R}^{3}$ be an open subset with coordinate $(x,y,z)$ and let $\pi_{U}:M_{U}\to U$ be a principal $S^{1}$-bundle. $V$ is a positive harmonic function defined on $U$ such that $\frac{1}{2\pi}\star dV$ represents the Chern class of this bundle, where $\star$ is the Hodge star operator with respect to the Euclidean metric $dg_{E}^{2}$ on $\mathbb{R}^{3}$. Then there is a connection $1$-form $\omega$ whose curvature is $\star dV$. The Gibbons-Hawking ansatz then produces a hyperkähler metric on $M_{U}$:
$g=\frac{1}{V}\omega^{2}+V\pi_{U}^{*}(dg_{E}^{2}).$
@@
$\Omega_{x}=dx\wedge\omega+Vdy\wedge dz,\ J_{x}:dx\mapsto V^{-1}\omega,\ dy\mapsto dz$
$\Omega_{y}=dy\wedge\omega+Vdz\wedge dx,\ J_{y}:dy\mapsto V^{-1}\omega,\ dz\mapsto dx$
$\Omega_{z}=dz\wedge\omega+Vdx\wedge dy,\ J_{z}:dz\mapsto V^{-1}\omega,\ dx\mapsto dy,$
@@
is a well-defined smooth harmonic function on $\mathbb{R}^{3}\setminus\{p_{j}\}$. One checks that the closed $2$-form $\frac{1}{2\pi}\star dV$ is a representative element of the Chern class $(-1,-1,\cdots)$ of the principal $S^{1}$-bundle. Hence there is a connection on $\pi_{0}$ with curvature $\star dV$. Let $\omega$ denote the connection form of this connection. The Gibbons-Hawking metric
$g=\frac{1}{V}\omega^{2}+V\pi_{0}^{*}(dg_{E}^{2})$
on $M_{0}$ extends smoothly across the added $S^{1}$-fixed points $\pi^{-1}(p_{j})\subset M$, and Anderson-Kronheimer-LeBrun [1] showed that the resulting metric $g$ on $M$ is complete and hyperkähler, while $M$ has infinite topological type.
In the sequel we shall mainly work with the complex structure $J:=J_{x}$. For later use it is convenient to describe it in slightly more concrete coordinates. Writing $\omega=dt+\theta$ for some $1$-form $\theta$ such that $d\theta=\star dV$, where $t$ is the angular coordinate along the $S^{1}$-fibres, the complex structure $J$ can be characterized by
$dx\mapsto V^{-1}(dt+\theta),$ (3)
$dy\mapsto dz.$ (4)
@@ 1.2 Main results
In this section, we present the main results derived from the hyperkähler manifolds of infinite topological type constructed in the previous section.
Throughout, let
  • $\{p_{j}\}_{j=1}^{\infty}\subset\mathbb{R}^{3}$ be a closed, discrete and countable set of points, with $p_{j}=(x_{j},y_{j},z_{j})$. We write $A=\{p_{j}:j\in\mathbb{N}^{+}\}$ and $U=\mathbb{R}^{3}\setminus A$.
  • $H:\mathbb{R}^{3}\to\mathbb{C}$ be the projection $H(x,y,z)=y+{\rm{i}}z$ and $a_{j}=H(p_{j})=y_{j}+{\rm{i}}z_{j}$.
We say that the configuration $\{p_{j}\}_{j=1}^{\infty}$ is generic if the projection
$H:\{p_{j}\}_{j=1}^{\infty}\to\mathbb{C}$
is injective, that is to say, $\{a_{j}\}_{j=1}^{\infty}$ are pairwise distinct.
@@
can be represented in the form of $\star dV_{\bm{e}}]$ where $V_{\bm{e}}$ is some positive harmonic function on $U$, if and only if $e_{j}\leq 0$ for all $j\in\mathbb{N}^{+}$ and there exists a point $x\in U$ such that
@@
In the concrete Gibbons-Hawking set-up in Section 1.1, the pricipal $S^{1}$-bundle $\pi_{0}:M_{0}\to\mathbb{R}^{3}\setminus\{p_{j}\}_{j=1}^{\infty}$ is chosen so that its first Chern class to be 1-1 on a small sphere enclosing each puncture $p_{j}$. Equivalently, the corresponding cohomology class $e=(e_{1},e_{2},\cdots)$ is fixed by the condition $e_{j}=-1$ for all $j\in\mathbb{N}^{+}$. Then we obtain a positive harmonic function $V_{\bm{e}}$ on $U=\mathbb{R}^{3}\setminus\{p_{j}\}_{j=1}^{\infty}$ of the form
@@
so that $\star dV_{\bm{e}}]=e$. This is exactly the potential appearing in (2) since $V=2\pi V_{\bm{e}}$. In particular, the requirement $e_{j}=-1$ for all $j$ ensures that the resulting hyperkähler metric extends smoothly across the added $S^{1}$-fixed points over $\{p_{j}\}_{j=1}^{\infty}$, producing the complete hyperkähler manifold $M$ of infinite topological type described in Section 1.1.
Our second main result describes $(M,J)$ under the generic hypothesis on the monopole points. For this we recall the standard Weierstrass primary factors
and define the entire function
$P(u)=u^{\delta}\prod_{a_{j}\neq 0}E_{j}\left(u/a_{j}\right)$
where $\delta=0$ if $a_{j}\neq 0$ for all $j$ and $\delta=1$ if $0\in\{a_{j}\}$.
Theorem 1.3 (Generic case).
Suppose that the puncture points $\{p_{j}\}$ satisfy
$\sum_{j=2}^{\infty}\frac{1}{\|p_{1}-p_{j}\|}<\infty,$
and that the configuration is generic in the sense that the projections $a_{j}=H(p_{j})$ are pairwise distinct. If, in addition, the set $\{a_{j}\}_{j=1}^{\infty}$ has no accumulation point on $\mathbb{C}$, then $(M,J)$ is biholomorphic to the hypersurface in $\mathbb{C}^{3}$ given by
$u_{1}u_{2}=P(u_{3}).$
Moreover, we could prove in Section 3 that for almost all complex structures in the $S^{2}$-family, $M$ is likewise biholomorphic to a hypersurface of this specific form in $\mathbb{C}^{3}$.
When the generic assumption fails, several punctures may project to the same point on $\mathbb{C}$ under $H$. In this non-generic situation $M$ can no longer be described directly as a hypersurface in $\mathbb{C}^{3}$; instead, it is realized as the minimal resolution of a singular hypersurface.
Theorem 1.4 (Non-generic case).
For countable distinct points $p_{j}\in\mathbb{R}^{3},\ j\in\mathbb{N^{+}}$ and
$\sum_{j=2}^{\infty}\frac{1}{\|p_{1}-p_{j}\|}<\infty,$
set $a_{j}=H(p_{j})=y_{j}+{\rm{i}}z_{j}$ and let $\{b_{1},b_{2},\cdots\}$ be the set of distinct values of $\{a_{j}\}$. Assume that
@@
    the set $\{b_{k}\}_{k=1}^{\infty}$ has no accumulation point on $\mathbb{C}$;
@@
    $m_{0}:=\#\{j\in\mathbb{N}^{+}:a_{j}=0\}<\infty,$
    $m_{k}:=\#\{j\in\mathbb{N^{+}}:a_{j}=b_{k}\}<\infty.$
Relabel the points $\{p_{j}\}_{j}$ as $\{p_{k,l}\}_{k,l}$ so that $H^{-1}(b_{k})=\{p_{k,l}:1\leq l\leq m_{k}\}$ for all $k\in\mathbb{N}^{+}$. Using suitable primary factors, we define an entire function
$P(u):=u^{m_{0}}\prod_{k:b_{k}\neq 0}\left(E_{c_{k}}\left(\frac{u}{b_{k}}\right)\right)^{m_{k}}$
where the constants $c_{k}$ are properly chosen so that the infinite product converge uniformly on every compact subset of $\mathbb{C}$.
Let $S$ in $\mathbb{C}^{3}$ be the surface defined by
$u_{1}u_{2}=P(u_{3}).$
Then $(M,J)$ is biholomorphic to the minimal resolution of $S$.
@@ 1.3 Outline
The paper is structured as follows. In Section 2, we investigate the Riesz measure of the harmonic function $V$ and give the proof of Theorem 1.1. In Section 3, we discuss the complex structure of manifold $M$ under a generic assumption. By constructing explicit local charts and analyzing the equivalence relation, we prove that $(M,J)$ is biholomorphic to a hypersurface in $\mathbb{C}^{3}$. Furthermore, we show in Corollary 3.7 that this hypersurface equivalence holds for almost all complex structures in the $S^{2}$-family. Finally, in Section 4, we relax the generic assumption and discuss the non-generic case. Under some specified conditions we prove that $(M,J)$ is the minimal resolution of a hypersurface, as stated in Theorem 1.4.
@@ 2 Existence criteria for harmonic functions via Riesz measures
In this section we justify the choice of Gibbons-Hawking potential $V$ and give the proof of Theorem 1.1. The strategy is to apply the uniqueness of the Riesz measure associated to a subharmonic function and to compute this measure explicitly in terms of the monopole configuration.
@@
Each $V_{\bm{e}}^{k}$ is a positive harmonic function on $U$, and the sequence $\{V_{\bm{e}}^{k}\}_{k\geq 1}$ is monotonically increasing. We invoke the following version of Harnack’s theorem (see [13, Theorem 1.20]).
@@
$\int_{\partial B(p_{j},\epsilon_{j})}\star dV_{\bm{e}}=e_{j}$
for $\epsilon_{j}$ small enough, which implies that $\star dV_{\bm{e}}]={\bm{e}}$.
@@
$[\star dV_{\bm{e}}]={\bm{e}}=(e_{1},e_{2},\cdots)\in H^{2}(U,\mathbb{Z}),$
@@
Therefore, if $V_{\bm{e}}$ is a positive harmonic function on $U$ such that $\star dV_{\bm{e}}]={\bm{e}}$, there exists a real harmonic function $V_{1}$ on $B(p_{1},\epsilon)$ and a real constant $b_{1}\geq 0$ such that
@@
$e_{1}=\int_{\partial B(p_{1},{\frac{\epsilon}{2}})}\star dV_{\bm{e}}=\int_{\partial B(p_{1},{\frac{\epsilon}{2}})}\star dV_{1}+\int_{\partial B(p_{1},{\frac{\epsilon}{2}})}\star d\left(b_{1}\cdot\frac{1}{4\pi|x-p_{1}|}\right)=-b_{1}.$
@@
To extract the additional summability condition, we now use the language of Riesz measures. We recall two standard results from the theory of subharmonic functions [13, Theorem 3.9, Theorem 3.20].
@@
$\int_{1}^{\infty}\frac{n(t)}{t^{m-1}}dt<\infty.$ (8)
@@
Assume $V_{\bm{e}}$ is a positive harmonic function on $U$ with $\star dV_{\bm{e}}]={\bm{e}}$. We extend it to a subharmonic function on $\mathbb{R}^{3}$ by setting
@@
In this way, the function $V_{\bm{e}}$ is upper semi-continuous on $\mathbb{R}^{3}$, satisfies $\infty\leq-V_{\bm{e}}(x)<\infty$, and obeys the mean-value inequality: for any point $x_{0}\in\mathbb{R}^{3}$ there exist a sufficiently small $r>0$ such that
$(-V_{\bm{e}})(x_{0})\leq\frac{1}{4\pi r^{2}}\int_{S(x_{0},r)}(-V_{\bm{e}})(x)d\sigma(x),$
where $d\sigma(x)$ denotes surface area on the sphere $S(x_{0},r)$. Define $u:=-V_{\bm{e}}$. Then $u$ is a subharmonic function on $\mathbb{R}^{3}$. Let $\mu$ be the Riesz measure of $u$ and set $n(t)=\mu(\overline{B(0,t)})$. We shall calculate $n(t)$.
@@
$\int_{B(p_{j},\epsilon_{j})}\varphi d\mu=-\int_{B(p_{j},\epsilon_{j})}\frac{|e_{j}|}{4\pi|x-p_{j}|}\cdot\Delta\varphi dx=|e_{j}|\cdot\varphi(p_{j}),$
@@
$\int_{1}^{\infty}\frac{n(t)}{t^{2}}dt$ $=\int_{1}^{|p_{N+1}|}\frac{1}{t^{2}}\left(\sum_{i=1}^{N}|e_{i}|\right)dt+\sum_{i\geq N+1}^{\infty}\int_{|p_{i}|}^{|p_{i+1}|}\frac{1}{t^{2}}\left(\sum_{j=1}^{i}|e_{j}|\right)dt$
@@
$\int_{1}^{\infty}\frac{n(t)}{t^{m-1}}dt$ $=\sum_{j=1}^{\infty}\frac{|e_{j}|}{|p_{j}|}.$

### v2

@@ Abstract
In this paper, we study the complex structures of complete hyperkähler four-manifolds of infinite topological type arising from the Gibbons-Hawking ansatz. We show that for almost all complex structures in the hyperkähler family, the manifold is biholomorphic to a hypersurface in $\mathbb{C}^{3}$ defined by an explicit entire function. For the remaining complex structures, we further prove that the manifold is biholomorphic to the minimal resolution of a singular surface in $\mathbb{C}^{3}$ under certain conditions. Thus, we partially extend LeBrun’s celebrated work [19] to the context of countably many punctures.
@@ 1 Introduction
In their pioneering work, Gibbons and Hawking [9, 10] introduced a construction of four-dimensional hyperkähler metrics with a tri-Hamiltonian $S^{1}$-action, now known as the Gibbons-Hawking ansatz. Hitchin [15], through the approach of twistor space, developed a coherent picture of the hyperkähler geometry behind these constructions. Kronheimer [18] described four-dimensional asymptotically locally Euclidean spaces arising as finite-dimensional hyperkähler quotients. Bielawski [3] analyzed complete hyperkähler manifolds with local tri-Hamiltonian actions and general dimension, placing a wide range of examples into a unified framework. In these works, the relevant configurations of puncture points are finite, and the resulting manifolds are of finite topological type. On the other hand, Anderson, Kronheimer and LeBrun [1] used Gibbons-Hawking metrics to construct complete Ricci-flat hyperkähler manifolds with infinite topological type. Goto [11] later constructed hyperkähler manifolds with infinite topological type and general dimension from the viewpoint of infinite dimensional hyperkähler quotients, closely related to Kronheimer’s construction [18] of hyperkähler manifolds of type $A_{k}$. Subsequent work by Hattori [12, 13] studied the volume growth and holomorphic symplectic structures of these type $A_{\infty}$ manifolds. Swann [20] provided a classification of complete four-dimensional hyperkähler manifolds that admit a tri-Hamiltonian $S^{1}$-action, without imposing any assumptions on their topological type. Dancer and Swann [6, 5] obtained further developments for general-dimensional hyperkähler manifolds of infinite topological type with a tri-Hamiltonian torus action.
In this manuscript, inspired by the works [1, 19, 17, 4], we investigate the complex structures of complete four-dimensional hyperkähler manifolds $M$ of infinite topological type obtained from the Gibbons-Hawking ansatz. Our first result (Theorem 1.1) gives an existence criterion for the harmonic functions in the Gibbons-Hawking ansatz for such manifolds. The corresponding existence statement was already proved in the work of Kalafat and Sawon [17] by an argument based on the maximum principle. We recover the criterion by a different method, using potential theory. Building on the method in LeBrun [19], we show that for almost all complex structures in the hyperkähler family on $M$, the underlying complex manifold is biholomorphic to a hypersurface in $\mathbb{C}^{3}$ defined by an explicit entire function (Theorem 1.3). LeBrun’s work treats the case where the circle action has finitely many fixed points and our description extends his picture to the countably many fixed points. For the remaining complex structures, which form a set of measure zero, we prove that under an additional finiteness condition, the complex manifold $M$ can be realized as the minimal resolution of a singular hypersurface in $\mathbb{C}^{3}$ (Theorem 1.4). In this way we relate the complex structures arising from the Gibbons–Hawking ansatz with infinite topological type directly to the geometry of the puncture set. We also note that Hattori [13] proved a similar classification theorem for four-dimensional hyperkähler manifolds of type $A_{\infty}$.
@@ 1.1 Background of Gibbons-Hawking ansatz
Let $U\subset\mathbb{R}^{3}$ be an open subset with coordinates $(x,y,z)$ and let $\pi_{U}:M_{U}\to U$ be a principal $S^{1}$-bundle. $V$ is a positive harmonic function defined on $U$ such that $\frac{1}{2\pi}\star{\mathrm{d}}V$ represents the Chern class of this bundle, where $\star$ is the Hodge star operator with respect to the Euclidean metric $g_{\text{Euc}}$ on $\mathbb{R}^{3}$. Then there is a connection $1$-form $\omega$ whose curvature is $\star{{\mathrm{d}}}V$. The Gibbons-Hawking ansatz then produces a hyperkähler metric on $M_{U}$:
$g=\frac{1}{V}\omega^{2}+V\pi_{U}^{*}g_{\text{Euc}}.$
@@
$\Omega_{x}={\mathrm{d}}x\wedge\omega+V{{\mathrm{d}}}y\wedge{{\mathrm{d}}}z,\ J_{x}:{{\mathrm{d}}}x\mapsto V^{-1}\omega,\ {{\mathrm{d}}}y\mapsto{{\mathrm{d}}}z$
$\Omega_{y}={{\mathrm{d}}}y\wedge\omega+V{{\mathrm{d}}}z\wedge{{\mathrm{d}}}x,\ J_{y}:{{\mathrm{d}}}y\mapsto V^{-1}\omega,\ {{\mathrm{d}}}z\mapsto{{\mathrm{d}}}x$
$\Omega_{z}={{\mathrm{d}}}z\wedge\omega+V{{\mathrm{d}}}x\wedge{{\mathrm{d}}}y,\ J_{z}:{{\mathrm{d}}}z\mapsto V^{-1}\omega,\ {{\mathrm{d}}}x\mapsto{{\mathrm{d}}}y,$
@@
is a well-defined smooth harmonic function on $\mathbb{R}^{3}\setminus\{p_{j}\}$. One checks that the closed $2$-form $\frac{1}{2\pi}\star{{\mathrm{d}}}V$ is a representative element of the Chern class $(-1,-1,\cdots)$ of the principal $S^{1}$-bundle. Hence there is a connection on $\pi_{0}$ with curvature $\star{{\mathrm{d}}}V$. Let $\omega$ denote the connection form of this connection. The Gibbons-Hawking metric
$g=\frac{1}{V}\omega^{2}+V\pi_{0}^{*}g_{\text{Euc}}$
on $M_{0}$ extends smoothly across the added $S^{1}$-fixed points $\pi^{-1}(p_{j})\subset M$, and Anderson-Kronheimer-LeBrun [1] showed that the resulting metric $g$ on $M$ is complete and hyperkähler, while $M$ has infinite topological type. For simplicity, we call this harmonic function $V$ the Gibbons-Hawking potential.
@@ 1.2 Main results
In this section, we present the main results derived from the hyperkähler manifolds of infinite topological type constructed in the previous subsection.
Throughout,
  • Let $\{p_{j}\}_{j=1}^{\infty}\subset\mathbb{R}^{3}$ be a closed, discrete and countable set of points, with $p_{j}=(x_{j},y_{j},z_{j})$. We take $A:=\{p_{j}:j\in\mathbb{N}^{+}\}$ and $U:=\mathbb{R}^{3}\setminus A$. Then
    $H^{2}(U,\mathbb{Z})=\prod_{j=1}^{\infty}\mathbb{Z}.$
  • Let $\mathbb{S}^{2}\subset\mathbb{R}^{3}$ be the unit sphere
    $\mathbb{S}^{2}:=\{(x,y,z)\in\mathbb{R}^{3}:x^{2}+y^{2}+z^{2}=1\}.$
    We denote the standard measure on $\mathbb{S}^{2}$ by $m$. For a direction $v\in\mathbb{S}^{2}$, we define the complex structure in the hyperkähler family of $M$ by
    $J_{v}:=v_{1}J_{x}+v_{2}J_{y}+v_{3}J_{z}.$
  • Let $\Pi_{v}$ be the orthogonal projection from $\mathbb{R}^{3}$ onto the plane $v^{\perp}$ orthogonal to $v$. Define $a_{j}(v):=\Pi_{v}(p_{j})$ and let $\{b_{1}(v),b_{2}(v),\cdots\}$ be the set of distinct values of $\{a_{j}(v)\}$.
@@
can be represented in the form of $\star{{\mathrm{d}}}V_{\bm{e}}]$ where $V_{\bm{e}}$ is some positive harmonic function on $U$, if and only if $e_{j}\leq 0$ for all $j\in\mathbb{N}^{+}$ and there exists a point $x\in U$ such that
@@
In the concrete Gibbons-Hawking set-up in Section 1.1, the principal $S^{1}$-bundle $\pi_{0}:M_{0}\to\mathbb{R}^{3}\setminus\{p_{j}\}_{j=1}^{\infty}$ is chosen so that its first Chern class takes the value 1-1 on a small sphere enclosing each puncture $p_{j}$. Equivalently, the corresponding cohomology class $e=(e_{1},e_{2},\cdots)$ is fixed by the condition $e_{j}=-1$ for all $j\in\mathbb{N}^{+}$. Then we obtain a positive harmonic function $V_{\bm{e}}$ on $U=\mathbb{R}^{3}\setminus\{p_{j}\}_{j=1}^{\infty}$ of the form
@@
so that $\star{{\mathrm{d}}}V_{\bm{e}}]=e$. This is exactly the harmonic function appearing in (2) since $V=2\pi V_{\bm{e}}$. In particular, the requirement $e_{j}=-1$ for all $j$ ensures that the resulting hyperkähler metric extends smoothly across the added $S^{1}$-fixed points over $\{p_{j}\}_{j=1}^{\infty}$, producing the complete hyperkähler manifold $M$ of infinite topological type described in Section 1.1.
Theorem 1.3.
Assume that the punctures $\{p_{j}\}$ satisfy
[…]
Then there exists a subset $B$ of $\mathbb{S}^{2}$ with $m(\mathbb{S}^{2}\setminus B)=0$ such that for every $v\in B$, $(M,J_{v})$ is biholomorphic to a hypersurface in $\mathbb{C}^{3}$ defined by
$u_{1}u_{2}=P_{v}(u_{3})$
where $P_{v}$ is an entire function determined by $\{p_{j}\}$ and $v$.
When $v\notin B$, $M$ can no longer be described directly as a hypersurface in $\mathbb{C}^{3}$. Nevertheless, under an additional finiteness assumption on the multiplicities of the projection points, $(M,J_{v})$ can be realized as the minimal resolution of a singular hypersurface.
Theorem 1.4.
Assume that the punctures $\{p_{j}\}$ satisfy
[…]
If $v\in\mathbb{S}^{2}\setminus B$ is a direction satisfies the following conditions:
@@
    The set $\{b_{k}(v)\}_{k=1}^{\infty}$ has no accumulation point on $\mathbb{C}$;
@@
    $m_{0}(v):=\#\{j\in\mathbb{N}^{+}:a_{j}(v)=0\}<\infty,$
    $m_{k}(v):=\#\{j\in\mathbb{N^{+}}:a_{j}(v)=b_{k}(v)\}<\infty.$
Then $(M,J_{v})$ is biholomorphic to the minimal resolution of a surface in $\mathbb{C}^{3}$ defined by
$u_{1}u_{2}=P_{v}(u_{3})$
where $P_{v}$ is an entire function determined by $\{p_{j}\}$ and $v$. In particular, the surface has at most countably many orbifold singularities of type $A_{m_{k}(v)-1}$ for each $m_{k(v)}$ greater than one.
@@ 1.3 Outline
The paper is structured as follows. In Section 2, we investigate the Riesz measure of the harmonic function $V$ and give the proof of Theorem 1.1. In Section 3, we first analyse the case for a fixed complex structure and prove our key Lemma 3.1, which gives the biholomorphism between $M$ and a hypersurface in $\mathbb{C}^{3}$. Combining this with a measure argument, we establish Theorem 1.3 for almost all complex structures in the hyperkähler family of $M$. Finally, in Section 4, we focus on the remaining complex structures and we prove that $M$ is the minimal resolution of a hypersurface under the additional finiteness condition, as stated in Theorem 1.4.
@@ 2 Existence criteria for harmonic functions via Riesz measures
In this section we justify the choice of Gibbons-Hawking potential $V$ and give the proof of Theorem 1.1. The strategy is to apply the uniqueness of the Riesz measure associated to a subharmonic function and to compute this measure explicitly in terms of the puncture configuration.
@@
Each $V_{\bm{e}}^{k}$ is a positive harmonic function on $U$, and the sequence $\{V_{\bm{e}}^{k}\}_{k\geq 1}$ is monotonically increasing. We invoke the following version of Harnack’s theorem (see [14, Theorem 1.20]).
@@
$\int_{\partial B(p_{j},\epsilon_{j})}\star{{\mathrm{d}}}V_{\bm{e}}=e_{j}$
for $\epsilon_{j}$ small enough, which implies that $\star{{\mathrm{d}}}V_{\bm{e}}]={\bm{e}}$.
@@
$[\star{{\mathrm{d}}}V_{\bm{e}}]={\bm{e}}=(e_{1},e_{2},\cdots)\in H^{2}(U,\mathbb{Z}),$
@@
Therefore, if $V_{\bm{e}}$ is a positive harmonic function on $U$ such that $\star{{\mathrm{d}}}V_{\bm{e}}]={\bm{e}}$, there exists a real harmonic function $V_{1}$ on $B(p_{1},\epsilon)$ and a real constant $b_{1}\geq 0$ such that
@@
$e_{1}=\int_{\partial B(p_{1},{\frac{\epsilon}{2}})}\star{{\mathrm{d}}}V_{\bm{e}}=\int_{\partial B(p_{1},{\frac{\epsilon}{2}})}\star{{\mathrm{d}}}V_{1}+\int_{\partial B(p_{1},{\frac{\epsilon}{2}})}\star{{\mathrm{d}}}\left(b_{1}\cdot\frac{1}{4\pi|x-p_{1}|}\right)=-b_{1}.$
@@
To extract the additional summability condition, we now use the language of Riesz measures. We recall two standard results from the theory of subharmonic functions [14, Theorem 3.9, Theorem 3.20].
@@
$\int_{1}^{\infty}\frac{n(t)}{t^{m-1}}{{\mathrm{d}}}t<\infty.$ (6)
@@
Assume $V_{\bm{e}}$ is a positive harmonic function on $U$ with $\star{{\mathrm{d}}}V_{\bm{e}}]={\bm{e}}$. We extend it to a subharmonic function on $\mathbb{R}^{3}$ by setting
@@
In this way, the function $V_{\bm{e}}$ is upper semi-continuous on $\mathbb{R}^{3}$, satisfies $\infty\leq-V_{\bm{e}}(x)<\infty$, and obeys the mean-value inequality: for any point $x_{0}\in\mathbb{R}^{3}$ there exists a sufficiently small $r>0$ such that
$(-V_{\bm{e}})(x_{0})\leq\frac{1}{4\pi r^{2}}\int_{S(x_{0},r)}(-V_{\bm{e}})(x){{\mathrm{d}}}\sigma(x),$
where ${{\mathrm{d}}}\sigma(x)$ denotes surface area on the sphere $S(x_{0},r)$. Define $u:=-V_{\bm{e}}$. Then $u$ is a subharmonic function on $\mathbb{R}^{3}$. Let $\mu$ be the Riesz measure of $u$ and set $n(t)=\mu(\overline{B(0,t)})$. We shall calculate $n(t)$.
@@
$\int_{B(p_{j},\epsilon_{j})}\varphi{{\mathrm{d}}}\mu=-\int_{B(p_{j},\epsilon_{j})}\frac{|e_{j}|}{4\pi|x-p_{j}|}\cdot\Delta\varphi{{\mathrm{d}}}x=|e_{j}|\cdot\varphi(p_{j}),$
@@
$\int_{1}^{\infty}\frac{n(t)}{t^{2}}{{\mathrm{d}}}t$ $=\int_{1}^{|p_{N+1}|}\frac{1}{t^{2}}\left(\sum_{i=1}^{N}|e_{i}|\right){{\mathrm{d}}}t+\sum_{i\geq N+1}^{\infty}\int_{|p_{i}|}^{|p_{i+1}|}\frac{1}{t^{2}}\left(\sum_{j=1}^{i}|e_{j}|\right){{\mathrm{d}}}t$
@@
$\int_{1}^{\infty}\frac{n(t)}{t^{m-1}}{{\mathrm{d}}}t$ $=\sum_{j=1}^{\infty}\frac{|e_{j}|}{|p_{j}|}.$